EVE AUG-4-99 shift summary

Alexander Ledovskoy (ledovsko@fnpxsr)
Thu, 5 Aug 1999 00:02:54 -0500

Shift Summary for EVE shift, 4-AUG-99

On Shift: S. Laplace, S. Ledovskoy

Hours of beam available at all intensities: 7.75h
Hours of beam available at nominal intensity: 5.92h
Hours of E832 data (special runs not included): 5.54h
Hours of special runs (laser scans, muon runs...):

Accounting of lost time: run stop'n'go, tape change. (0.38h)

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:

QKA006
QKCC09
QKCR05
QKLA11
QKMU11
QKNO18
QKNL17
QKPE09
QKSM09
QKB074-075
QKC210-215
QKK210-215
QKS210-215

========================================================================
Accesses during shift: none
========================================================================
Overview of shift: E823 data taking at ~6.8E+12 Beam went away last 45min.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14105 146 15:30 New shift crew takes over smooth data taking.

151 ktevana errors, KFREND fixes it.

16:00 SEM is 7.15E12 last couple of spills. Call MCR.

170 ktevana errors, KFREND fixes it.

17:00 Sandrine checked onlinee histos:
CsI: 1497 shows RMS 3.1% for several spills
DC: Track link efficiency vs position inside cell -
center of "beam region" curve is under the limit
Ch1 multiplicity has peak at 4.79
KQ: unpack errors are seen in YTF plots,
CH3Y has missing KQ ORs
However, e8323#kqke3 looks OK

18:11 14105 ends gracefully with tapes full.
Change tapes, initialize new run.

14106 18:40 New run started. Online histos from previous run
disappeared!

18 CsI 1393 has RMS 36% Back to normal next spill.

56 ktevana errors, 2 KFRENDs fix it.

58 LOTs of CsI channels have RMS above 10%
Back to normal next spill

18:50 3 new Ar-Ethane bottles now online west side.

114 ktevana errors, KFREND fixes it.

21:15 Ed called. He modified prescale in destination logic
for Ke3 from 30 to 50 because we are running at
higher intensity now. This change will be in
effect starting next run

172 ktevana errors, KFREND fixes it.

23:16 Beam goes away. MI kicker PS problems. Estimate 20min

23:44 New estimate for beam is 1.5 hours

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift: Keep taking data.